# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v1

@@ 3.1. Entities
The IAI Model comprises six core entities: Human (H), Interaction (I), Artifact (A), Text Prompt (T), Augmented Instruction (Aug), and Generative AI (G). Concretely, we treat an entity as a semantically coherent object or agent in the interaction ecology. Table 1 provides an overview for all entities and relations between them.
@@
First, to capture the qualitative differences between modalities, we separate Text Prompt (T) from Interaction (I). Conceptually, T is a free-form, general-purpose, natural language specification of intent (e.g., make the flower red”). I denotes focused, modality-specific operations (e.g., click, brush, drag, sketch, widget selection) that supply concrete referents or constraints (e.g., a brush mask, a bounding box, a selected code block). This separation matters because T and I differ sharply in expressivity, precision, and in how the model understands them. However, such separation is not sufficient to represent what the model actually consumes in many practical GenAI workflows.
@@
To address the issue, secondly, we introduce Augmented Instruction (Aug) as a new entity. In GenAI system interfaces, text prompts and interactions are not independent “inputs” that the model somehow interprets in isolation. Interactions often produce structured, non-linguistic constraints (e.g., pixel masks, coordinate ranges, AST node identifiers, selected table rows or text segments, slider parameters) that must be encoded, normalized, and attached to a prompt in a machine-readable form (Shen et al., 2024b; Chung and Adar, 2023; Liu et al., 2025, 2023b; Almeda et al., 2024). Viewing Aug as the single, explicit input to GenAI simplifies paradigm comparisons, as tools differ mainly in how they combine T, I, and A to build Aug.
@@
Third, during iteration we consolidated several auxiliary entities (found in Sec. 2.2) into the six core entities to keep the model parsimonious while maintaining expressiveness. Concretely: (1) Context inputs (e.g., retrieved passages, prior conversation state, uploaded reference files) are modeled as part of Artifact (A) because they function as domain objects that ground instructions (Gao et al., 2024). Their inclusion is modeled via A $\rightarrow$ Aug (artifact-derived constraints) or A $\rightarrow$ G (artifact supplied as raw model context). (2) Temporary interaction products (e.g., highlighted spans, sketched regions, intermediate masks, or annotation buffers) are treated as manifestations of Interaction (I) rather than independent entities (Hu et al., 2025). They have only transient semantic life: they exist to constrain or point to artifacts or prompts and then are encoded into instructions or discarded. (3) Widgets have two aspects: when a widget is a UI primitive invoked by the human it is part of Interaction (H $\rightarrow$ I) (Feng et al., 2024; Masson et al., 2024), and when it is generated by the model it is represented as an AI $\rightarrow$ interaction proposal (G $\rightarrow$ I) that the human may accept (I $\rightarrow$ T (Andrews et al., 2024), I $\rightarrow$ Aug (Wang et al., 2024a), or I $\rightarrow$ A (Vaithilingam et al., 2024)). This treatment preserves provenance without inflating the entity set. (4) External tool invocation is treated as part of GenAI’s internal behavior (G $\rightarrow$ A). Many systems route domain-specific executors (e.g., image editors, compilers, specialized APIs) behind the model; these are execution mechanisms rather than affecting interaction paradigm design (Shen et al., 2025b). Modeling them as separate entities would conflate execution architecture with the interaction paradigms we aim to capture.
@@
Putting these together yields six entities (Table 1), and each plays a distinct semantic role. Conceptually the six-entity set is necessary: removing any entity collapses an entire class of workflows, e.g., without I, interaction-augmented cases degenerate to prompt-only; without T, prompt-only systems vanish; without Aug, interaction-infused instructions cannot be distinguished from raw prompts; without A, there is no target object; without G or H, agency is undefined. They are also sufficient to express human–GenAI communication paradigms with the interplay of interactions and prompts, as they are a universe of all nodes in both prompt-driven interaction flow and GUI interaction flow. This point is further justified through revisiting existing research tools in Sec. 4.
@@
Human (H) The end user who expresses intent and interacts with the GenAI system interface via text prompts and GUI actions. Cannot manipulate the AI directly except through these inputs. H $\rightarrow$ T: The user writes and refines a text prompt, which is the primary natural-language instruction conveying intent.
H $\rightarrow$ I: The user performs interactive actions (e.g., clicking, highlighting, dragging) on the interface or artifacts to supplement or refine the prompt.
Text Prompts (T) Natural-language instructions written by the user to convey intent. Intuitive but may be ambiguous or incomplete. Used alone or as part of a richer instruction for GenAI T $\rightarrow$ Aug: The text prompt is incorporated into the augmented instruction (combined with interaction-derived information).
T $\rightarrow$ G: The text prompt alone is sent to the GenAI as input for generating or operating on artifacts.
Interaction (I) Supplemental user actions (e.g., clicking, selecting, annotating) that add constraints or context to prompts. They refine the AI’s understanding but do not generate output independently. I $\rightarrow$ Aug: Interaction inputs are integrated into the augmented instruction, adding detail or constraints to the original prompt.
I $\rightarrow$ T: An interaction may modify the text prompt itself.
I $\rightarrow$ A: The user’s interactions act directly on domain artifacts (e.g., clicking/highlighting parts of an image or document) to specify or restrict the scope of the GenAI task.
Augmented Instruction (Aug) The combined instruction delivered to GenAI, formed by merging the text prompt with information derived from interactions. It encodes the complete intent for the AI and only exists as an input to the GenAI system. Aug $\rightarrow$ G: The augmented instruction is passed to the GenAI for execution. GenAI uses this enriched instruction to generate content or perform actions.
Artifacts (A) Domain objects (e.g., text, image, code) that GenAI operates on. They are the targets of user interactions and AI outputs, but not instructions themselves. A $\rightarrow$ Aug: User interactions on an artifact (e.g., highlighting a paragraph) are incorporated into augmented instruction.
A $\rightarrow$ G: Artifacts or extracted features used as contextual input to GenAI.
Generative AI (G) The model (e.g., LLM or diffusion model) that interprets the augmented instruction and generates or edits artifacts. It acts only upon receiving input. G $\rightarrow$ A: Upon receiving the augmented instruction, GenAI operates on artifacts by generating new content or triggering operations on domain objects (e.g., creating an image, editing a document).
G $\rightarrow$ I: Can initiate interactions (e.g., suggesting follow-up options or UI elements for user action).
@@ 3.2. Relations
Taken together, these principles justify modeling only a small, purposeful subset of all possible relations rather than the full combinatorial space. For each entity pair (X, Y), we evaluate whether X can produce information or control that Y can meaningfully consume in the context of human–GenAI communication. If so, the relation is included with its semantics recorded; if not, it is excluded with an explicit rationale. Table 1 presents the resulting relation set. Detailed explanations are as follows:
@@
  • Text Prompt (T). T is a general-purpose, free-form specification of intent. It can be sent directly to GenAI (T $\rightarrow$ G) or be combined with interaction-derived information to form an augmented instruction (T $\rightarrow$ Aug). T does not act on artifacts directly (no T $\rightarrow$ A).
  • Interaction (I). Interactions are focused, often single-purpose operations. Interactions can operate on artifacts (I $\rightarrow$ A) like selecting target elements, modifying or augmenting prompts (I $\rightarrow$ T), or feeding information into the augmented instruction (I $\rightarrow$ Aug). Interactions cannot evoke GenAI to generate by themselves (no I $\rightarrow$ G): they are not generators but mediators of specificity.
  • Augmented Instruction (Aug). Augmented instruction represents the instructions beyond pure NL prompts that the GenAI will execute. By definition, Aug is constructed from prompt and interaction inputs (T $\rightarrow$ Aug, I $\rightarrow$ Aug) and can additionally incorporate direct artifact-derived context (A $\rightarrow$ Aug) when a selection or context snippet is encoded into the instruction. The only valid execution path from Aug is into the model (Aug $\rightarrow$ G); Aug does not itself perform artifact edits (no Aug $\rightarrow$ A). Treating Aug as the single, explicit input to GenAI makes paradigm comparisons straightforward: different tools differ chiefly in how they build Aug (which combinations of T, I, and A feed into it). Please refer to Sec. 4.2 for more details.
  • Artifact (A). Artifacts are the domain objects—texts, images, code, datasets—that provide both targets and context. Artifacts are passive in the relation set: they do not autonomously produce text prompts (no A $\rightarrow$ T) or initiate interactions (no A $\rightarrow$ I). Relevant relations include I $\rightarrow$ A (interactions operate on artifacts to select or annotate), A $\rightarrow$ Aug (artifact content or references can be incorporated into the augmented instruction), and A $\rightarrow$ G (artifacts can be provided directly as model input). The distinction between A $\rightarrow$ Aug and A $\rightarrow$ G is meaningful for paradigm design: A $\rightarrow$ Aug indicates artifact-derived constraints become part of the composite instruction, whereas A $\rightarrow$ G indicates the artifact (as well as its extracted features) is supplied as raw model context.
@@ 3.3. Atomic Paradigm Graph
Fig. 2 depicts the full interaction-augmented instruction model as a directed graph. However, considering the task-specific needs and the complexity of integrating all entities and relations, not all of them must be adopted to represent an individual interface. Therefore, we introduce the concept of an atomic paradigm graph, defined as a minimal, self-contained subgraph of the model that captures a single, coherent interaction paradigm (Table 2). Here, atomic means that GenAI (G) plays only one role within the paradigm (avoiding overlapping or conflated functions) and that the paradigm necessarily involves interaction (I), as paradigms without interaction fall outside our scope. Each atomic paradigm graph is constructed by sequentially selecting entities and relations in the order they are enacted, thereby encoding the workflow through which a tool supports human–GenAI communication. Accordingly, a concrete tool can be represented by one or more atomic paradigm graphs, depending on how many distinct paradigms it supports. For example, an advanced image editing interface with brush-based region selection (Table 2-P4) would add H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug, and route the prompt through T $\rightarrow$ Aug $\rightarrow$ G, forming a richer paradigm (Masson et al., 2024; Liu et al., 2025). In another variant (Table 2-P8), if GenAI generates candidate widgets to further tune generated artifacts, the graph will include G $\rightarrow$ I (GenAI initiates widget interactions) followed by H $\rightarrow$ I $\rightarrow$ A (human interacts with widgets to act on artifacts) (Vaithilingam et al., 2024). These small structural differences capture the paradigm characteristics.
Table 2. Taxonomy of atomic interaction paradigms in human-GenAI communication.
Interaction Timing Starting
Resources
Paradigm Name Paradigm Description Atomic
Paradigm Graph
Related
Tools
Interaction Before Calling GenAI Prompt-only (no artifact at hand) P1. Interactive Prompt Enhancement Human selects parts of drafted prompts for GenAI to refine or expand content. [Uncaptioned image] (Almeda et al., 2024; Brade et al., 2023)
P2. Interactive Prompt Organization Human organize multiple text prompts into structured formats (e.g., tree) to fit the application. [Uncaptioned image] (Feng et al., 2024; Yen et al., 2024; Wu et al., 2022; Gmeiner et al., 2025; Nguyen et al., 2025)
P3. Interaction as Instruction Human’s interactions outside artifacts are included with prompts for GenAI to operate on artifacts. [Uncaptioned image] (Lin et al., 2025; Kim et al., 2023a; Zhou et al., 2025)
Artifact-grounded P4. Artifact as Instruction Human directly manipulates artifacts, which are combined with prompts as GenAI instructions. [Uncaptioned image] (Masson et al., 2024; Wang et al., 2024c; Chen et al., 2024; Fok et al., 2024; Liu et al., 2025, 2023a)
Interaction After Calling GenAI Prompt-only (no artifact at hand) P5. AI-driven Prompt Suggestion GenAI suggests new or extended prompts from the human’s initial input for human selection. [Uncaptioned image] (Bursztyn et al., 2021; Suh et al., 2023; Baek et al., 2023; Andrews et al., 2024; Angert et al., 2023; Wang et al., 2024a; Ma et al., 2024; Gmeiner et al., 2025)
P6. AI-driven Prompt Decomposition GenAI restructures prompts into fine-grained, organized forms for interactive manipulation. [Uncaptioned image] (Cai et al., 2024; Zhang et al., 2025)
P7. Generative Prompt Control Widgets GenAI generates interactive widgets for fine-grained prompt control. [Uncaptioned image] (Wang et al., 2024a; Gmeiner et al., 2025; Chen et al., 2025a)
P8. Generative Artifact Control Widgets GenAI generates widgets for humans to further manipulate or confirm artifact-related actions. [Uncaptioned image] (Vaithilingam et al., 2024; Suh et al., 2024; Gero et al., 2022; Petridis et al., 2024; Chen et al., 2024; Jiang et al., 2023; Wu et al., 2023)
Artifact-grounded P9. Artifact to Organized Instruction GenAI uses artifacts as starting points to generate structured prompts. [Uncaptioned image] (Zhang et al., 2023; Kim et al., 2023b; Gmeiner et al., 2025)
P10. Artifact to Multimodal Instruction GenAI parses and integrates artifacts to construct instructions. [Uncaptioned image] (Singh et al., 2024)
P11. Artifact-driven Prompt Enhancement GenAI suggests actions based on contextual requests integrating artifacts and prompts. [Uncaptioned image] (Chen et al., 2023)
P12. Interactive Artifact Refinement GenAI analyzes artifacts and initiates interactions based on prompts. [Uncaptioned image] (Tang et al., 2024)
@@ 4. Design Paradigms
4.1. Revisit the Corpus
Data. We revisited prior corpus about interaction-enhanced GenAI interfaces (Shen et al., 2025b; Gao et al., 2024; Hu et al., 2025; Luera et al., 2024; Lehmann and Buschek, 2024) and filtered tools according to three criteria: (1) the system involves at least one GenAI model; (2) text prompts are supported as a communication channel; and (3) at least one other interaction modality (e.g., selection, brushing, sketching) is used to augment text prompts. Applying these filters yielded 66 representative system interfaces for analysis [^1] The complete annotation results: https://interaction-augmented-instruction.github.io/.
Annotation and Analysis. Following Sec. 3.3, we decomposed each system interface into one or more atomic paradigm graphs (Fig. 2) through manual annotation. During annotation, each atomic graph is instantiated by choosing the entities and relations the system interface implements, and by assigning sequence indices to relations to indicate the temporal or information-flow order. Concurrent flows receive the same index (e.g., when a user selection of artifact elements is incorporated into an augmented instruction simultaneously: H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug). Two authors jointly performed the annotation: each coded half the corpus, cross-checked the other’s work, and resolved disagreements through iterative discussion until consensus was reached. Grounded in IAI model design (Sec. 3) and definitions (Table 1), our annotation followed these principles:

### v2

@@ 3.1. Entities
The IAI Model comprises six core entities: Human (H), Interaction (I), Artifact (A), Text Prompt (T), Augmented Instruction (Aug), and Generative AI (G). Concretely, we treat an entity as a semantically coherent object or agent in the interaction ecology. Tab. 1 provides an overview for all entities and relations.
@@
First, to capture the qualitative differences between modalities, we separate Text Prompt (T) from Interaction (I).   Conceptually, T is a free-form, general-purpose, natural language specification of intent (e.g., a stylized portrait of a young woman” in Fig. 2-a). I denotes focused, non-textual, often GUI user actions (e.g., click in Fig. 2-b, sketch in Fig. 2-c, brush in Fig. 2-d, drag in Fig. 2-e, widget selection in Fig. 2-f) that can modify, supplement, or operationalize concrete referents or constraints (e.g., a brush mask, a bounding box, a selected code block). This separation matters because T and I differ sharply in expressivity, precision, and in how the model understands them. However, such separation is not sufficient to represent what the model actually consumes in many practical GenAI workflows.
@@
To address the issue, secondly, we introduce Augmented Instruction (Aug) as a new entity. In GenAI system interfaces, text prompts and interactions are not independent “inputs” that the model somehow interprets in isolation. Interactions often produce structured, non-linguistic constraints (e.g., pixel masks, coordinate ranges, AST node identifiers, selected table rows or text segments) that must be encoded, normalized, and attached to a prompt in a machine-readable form (Shen et al., 2024a; Chung and Adar, 2023; Liu et al., 2025, 2023a; Almeda et al., 2024).   Viewing Aug as the explicit input to GenAI simplifies paradigm comparisons, as tools differ mainly in how they combine T, I, and A to build Aug.
@@
Third, during iteration we consolidated several auxiliary entities (found in Sec. 2.2) into the six core entities to keep the model parsimonious while maintaining expressiveness. Concretely: (1) Context inputs (e.g., retrieved passages, prior conversation state, uploaded reference files) are modeled as part of Artifact (A) because they function as domain objects that ground instructions (Gao et al., 2024). Their inclusion is modeled via A $\rightarrow$ Aug (artifact-derived constraints) or A $\rightarrow$ G (artifact supplied as raw model context). (2) Temporary interaction products (e.g., highlighted spans, sketched regions, intermediate masks, or annotation buffers) are treated as manifestations of Interaction (I) rather than independent entities (Hu et al., 2025). They have only transient semantic life: they exist to constrain or point to artifacts or prompts and then are encoded into instructions or discarded. (3) Widgets have two aspects: when a widget is a UI primitive invoked by the human it is part of Interaction (H $\rightarrow$ I) (Feng et al., 2024; Masson et al., 2024), and when it is generated by the model it is represented as an AI $\rightarrow$ interaction proposal (G $\rightarrow$ I) that the human may accept (I $\rightarrow$ T (Andrews et al., 2024), I $\rightarrow$ Aug (Wang et al., 2024d), or I $\rightarrow$ A (Vaithilingam et al., 2024)). This treatment preserves provenance without inflating the entity set. (4) External tool invocation is treated as part of GenAI’s internal behavior (G $\rightarrow$ A). Many systems route domain-specific executors (e.g., image editors, compilers, specialized APIs) behind the model; these are execution mechanisms rather than affecting interaction paradigm design (Shen et al., 2025a). Modeling them as separate entities would conflate execution architecture with the interaction paradigms we aim to capture.
@@
Putting these together yields six entities (Tab. 1), and each plays a distinct semantic role. Conceptually the six-entity set is necessary: removing any entity collapses an entire class of workflows, e.g., without I, interaction-augmented cases degenerate to prompt-only; without T, prompt-only systems vanish; without Aug, interaction-infused instructions cannot be distinguished from raw prompts; without A, there is no target object; without G or H, agency is undefined. They are also sufficient to express human–GenAI communication paradigms with the interplay of interactions and prompts, as they are a universe of all nodes in both prompt-driven interaction flow and GUI interaction flow. This point is further justified through revisiting existing research tools in Sec. 4.
@@
Human (H) The end user who holds intent and externalizes it through text prompts and interactions. Humans do not act on GenAI directly, but influence its behavior only via mediated instructions. H $\rightarrow$ T: The user authors or revises a text prompt to express open-ended goals in natural language.
H $\rightarrow$ I: The user performs interactive actions (e.g., clicking, highlighting, dragging) on the interface or artifacts to provide information.
Text Prompts (T) Natural language instructions written by the user to convey intent. Intuitive but may be ambiguous or incomplete. Used alone or as part of a richer instruction for GenAI T $\rightarrow$ Aug: Text prompts are incorporated into the augmented instruction (combined with interaction-derived information or artifacts).
T $\rightarrow$ G: The text prompt alone is directly submitted to GenAI as a complete instruction.
Interaction (I) Supplemental user actions (e.g., clicking, selecting, annotating) that concretize or constrain intent. Interactions do not generate content; they specify how or where intent applies. I $\rightarrow$ Aug: Interaction-derived information is integrated into the augmented instruction.
I $\rightarrow$ T: Interactions may modify the text prompt itself.
I $\rightarrow$ A: The user’s interactions operate on artifacts (e.g., highlighting parts of an image) to specify or restrict the target scope.
Augmented Instruction (Aug) The combined instruction delivered to GenAI, formed by merging information derived from text prompts, interactions, and artifacts. It encodes the complete intent for the AI and only exists as an input to the GenAI system. Aug $\rightarrow$ G: The augmented instruction is passed to GenAI for execution. GenAI uses this enriched instruction to generate content or perform actions.
Artifacts (A) Domain objects (e.g., text, image, code) that serve as both the state users interact with and the possible results of GenAI execution. Artifacts are not instructions. A $\rightarrow$ Aug: User interactions on artifacts (e.g., highlighting a paragraph) are incorporated into augmented instruction.
A $\rightarrow$ G: Artifacts provide contextual input or constraints for GenAI processing.
Generative AI (G) The model (e.g., LLM, diffusion model) that interprets instructions and operate on artifacts. Depending on the paradigm, GenAI may act as different roles (e.g., generator or clarifier). G $\rightarrow$ A: GenAI generates, transforms, or edits artifacts based on the instruction (e.g., creating an image, editing a document).
G $\rightarrow$ I: GenAI initiates interactions (e.g., suggesting follow-up options or UI elements) to solicit further user input.
@@ 3.2. Relations
Taken together, these principles justify modeling only a small, purposeful subset of all possible relations rather than the full combinatorial space. For each entity pair (X, Y), we evaluate whether X can produce information or control that Y can meaningfully consume in the context of human–GenAI communication. If so, the relation is included with its semantics recorded; if not, it is excluded with an explicit rationale. Tab. 1 presents the resulting relation set. Detailed explanations are as follows:
@@
  • Text Prompt (T).   T is a general-purpose, free-form specification of intent in a natural language string, which is manually authored or edited by a human. As an entity, T represents the text that will be directly sent to GenAI (T $\rightarrow$ G) or be combined with interaction-derived information to form an augmented instruction (T $\rightarrow$ Aug). T does not act on artifacts directly (no T $\rightarrow$ A), and only textual content counts as T; GUI actions or structured editing widgets are not defined as prompts.
  • Interaction (I).   Interaction (I) refers to non-textual, often GUI user actions (e.g., clicking, dragging, brushing, or manipulating structured editors) that can modify, supplement, or operationalize what the text prompt expresses or what the artifact provides, while not themselves constituting valid GenAI inputs. Interactions can operate on artifacts (I $\rightarrow$ A) like selecting target elements, modifying or augmenting prompts (I $\rightarrow$ T), or feeding information into the augmented instruction (I $\rightarrow$ Aug). Interactions cannot evoke GenAI to generate by themselves (no I $\rightarrow$ G): they are not generators but mediators of specificity. Furthermore, I excludes the act of writing text prompts as natural language strings without formatting.
  • Augmented Instruction (Aug). Augmented instruction represents the instructions beyond pure NL prompts that the GenAI will execute. By definition, Aug is constructed from prompt and interaction inputs (T $\rightarrow$ Aug, I $\rightarrow$ Aug) and can additionally incorporate direct artifact-derived context (A $\rightarrow$ Aug) when a selection or context snippet is encoded into the instruction. The only valid execution path from Aug is into the model (Aug $\rightarrow$ G); Aug does not itself perform artifact edits (no Aug $\rightarrow$ A).   Treating Aug as the explicit input to GenAI makes paradigm comparisons straightforward: different tools differ chiefly in how they build Aug (which combinations of T, I, and A feed into it). Please refer to Sec. 4.2 for more details. Meanwhile, T $\rightarrow$ G still remains essential, as users often craft a text prompt before the model initiates further interaction.
  • Artifact (A). Artifacts are the domain objects—texts, images, code, datasets—that provide both targets and context. Artifacts are passive in the relation set: they do not autonomously produce text prompts (no A $\rightarrow$ T) or initiate interactions (no A $\rightarrow$ I). Relevant relations include I $\rightarrow$ A (interactions operate on artifacts to select or annotate), A $\rightarrow$ Aug (artifact content or references can be incorporated into the augmented instruction), and A $\rightarrow$ G (artifacts can be provided directly as model input). The distinction between A $\rightarrow$ Aug and A $\rightarrow$ G is meaningful for paradigm design: A $\rightarrow$ Aug indicates artifact-derived constraints become part of the composite instruction, whereas A $\rightarrow$ G indicates the artifact (and its extracted features) is supplied as raw model context.
@@ 3.3. Atomic Paradigm Graph
Fig. 3 depicts the full interaction-augmented instruction model as a directed graph. However, considering the task-specific needs and the complexity of integrating all entities and relations, not all of them must be adopted to represent an individual interface. Therefore, we introduce the concept of an atomic paradigm graph, defined as a minimal, self-contained subgraph of the model that captures a single, coherent interaction paradigm (Tab. 2). Here, atomic means that GenAI (G) plays only one role within the paradigm (avoiding overlapping or conflated functions) and that the paradigm necessarily involves interaction (I), as paradigms without interaction fall outside our scope. Each atomic paradigm graph is constructed by sequentially selecting entities and relations in the order they are enacted, thereby encoding the workflow through which a tool supports human–GenAI communication. Accordingly, a concrete tool can be represented by one or more atomic paradigm graphs, depending on how many distinct paradigms it supports. For example, an advanced image editing interface with brush-based region selection (Tab. 2-P4) would add H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug, and route the prompt through T $\rightarrow$ Aug $\rightarrow$ G, forming a richer paradigm (Masson et al., 2024; Liu et al., 2025). In another variant (Tab. 2-P8), if GenAI generates widgets to further tune generated artifacts, the graph will include G $\rightarrow$ I (GenAI initiates widget interactions) followed by H $\rightarrow$ I $\rightarrow$ A (human interacts with widgets to act on artifacts) (Vaithilingam et al., 2024). These small structural differences capture the paradigm characteristics.
@@ 4. Design Paradigms
4.1. Revisit Corpora
Data. We revisited prior corpora about interaction-enhanced GenAI interfaces (Shen et al., 2025a; Gao et al., 2024; Hu et al., 2025; Luera et al., 2024; Lehmann and Buschek, 2024), which together enumerate 483 candidate instances, and filtered tools according to three criteria: (1) the system involves at least one GenAI model; (2) text prompts are supported for communication; and (3) at least one other interaction modality (e.g., selection, brushing, sketching) is used to augment text prompts. For each candidate, we examined available demo videos and project pages; otherwise we carefully read the paper to verify compliance. Three authors conducted the initial screening in roughly equal partitions, rotating assignments to ensure every decision was cross-checked. Applying these filters yielded 66 representative system interfaces for analysis.
Annotation and Analysis. Following Sec. 3.3, we decomposed each system interface into one or more atomic paradigm graphs (Fig. 3) through manual annotation. During annotation, each atomic graph is instantiated by choosing the entities and relations the system interface implements, and by assigning sequence indices to relations to indicate the temporal or information-flow order. Concurrent flows receive the same index (e.g., when a user selection of artifact elements is incorporated into an augmented instruction simultaneously: H $\rightarrow$ I $\rightarrow$ A $\rightarrow$ Aug).
Annotation was conducted by two authors using a structured, multi-stage protocol to ensure reliability and reproducibility. We began with a calibration pilot of 10 randomly sampled systems to test the initial codebook (grounded in IAI model design (Sec. 3) and definitions (Tab. 1), surface ambiguous cases, and refine operational rules. The two annotators then independently coded each system and cross-checked every annotation to reach a shared interpretation grounded in the codebook. All remaining uncertainties were reviewed and resolved with the involvement of a third author, and the resulting consensus annotations form the basis of our analysis. In addition to basic model definitions, our annotation followed these principles: